# Ordering the topological order in the fractional quantum Hall effect

### v2

@@ 8.1 Quantum dimension
Alternatively, the original motivation for this definition came from the underlying RCFT, where it is related to the dimension of a representation $a$ of a chiral algebra. Since these representations are infinite-dimensional, they do not have a standard notion of dimension. However, we can use the characters of these representations $\chi_{a}(\tau)$ (with $\tau$ the modular parameter), to define the quantum dimension of $a$ as [102]
@@
8.2 Single layer spinc theories
@@
Most historical examples of odd denominator FQH states have the order of the vison equal to $q$, i.e., $\ell=1$. Thus, in this section, we will take $\ell=1$, but generalize to larger values in the next section. Here, $s=q$ and $r=p$ and hence, the symmetry lines are ${\cal V}^{q,p}$. As mentioned in Sec. 7, this is a complete TQFT and the full TQFT thus factorizes as ${\cal V}^{q,p}\boxtimes{\cal T^{\prime}}$, where anyons in ${\cal T^{\prime}}$ carry integer charges. In this case, one can attach electrons to the $\mathcal{T}^{\prime}$ anyons to make all of them charge-neutral, so the theory can be formed with only bosonic degrees of freedom, i.e., a purely bosonic TO. This is formalized in Thm. C.17 in App. C.
Let us first discuss a few examples with trivial $\cal T^{\prime}$. These examples were the first FQH states to be described. Starting with Laughlin [41], wave-function-based approaches to FQH states were developed [106, 107, 108, 109, 110]. In the Laughlin states at filling $\nu=1/q$, the TQFT is just our $\mathcal{V}^{q,1}$ theory. In Chern-Simons language, these are the $\mathrm{U}(1)_{q}$ theories.
Now we will show how more general Abelian FQH states fit in. First, hierarchical FQH states can be represented by a $K$-matrix that is tri-diagonal with ones above and below the diagonal [111] and a $t$ vector given by $t_{ij}=\delta_{i1}$. To satisfy the spinc condition, we need to have $K_{11}\in 2\mathbb{Z}+1$ and $K_{jj}\in 2\mathbb{Z}$ for $j>1$. These theories are just the $\mathcal{V}^{q,p}$ minimal theories with $q=|\det K|$ and $p=qK^{-1}_{1,1}$. They are generated by the vison.
@@
These states have filling $\nu=n/(2np\pm 1)$, and a $\mathbb{Z}_{2np\pm 1}^{(1)}$ one-form symmetry group. As before, the vison generates the entire symmetry group, and so the Jain states are really just the $\mathcal{V}^{2np\pm 1,n}$ minimal theories. In fact, all the states that arise from additional flux attachments to these Jain states are the $\mathcal{V}^{q,p}$ minimal models, as shown in App. A of [88].
@@
Some examples with non-trivial $\cal T^{\prime}$ are the Read-Rezayi states at filling $\nu=p/q$ with $p$ and $q$ odd. Let us review this fact. The Read-Rezayi states have filling fractions $\nu=k/(km+2)$ for $k,m\in\mathbb{Z}$ [114, 98]. For odd $m$, these are spinc theories, while for even $m$ they are bosonic. In all cases, they are constructed as the following Chern-Simons theory [98, 115]
@@
Here, the theory ${\cal F}_{k}$ is defined in App. A.3. Thus, these states are all minimal states stacked with an additional theory. For $\nu=13/5=2+3/5$ this additional theory is ${\cal F}_{3}$, the Fibonacci anyon model. The case of $\nu=12/5=2+2/5$ is the particle-hole conjugate of $13/5$. These states have been proposed to be realized at these fillings [116, 117, 118, 119, 120], though there is some debate over their applicability.
@@
where the subscript indicates that we are gauging the $\mathbb{Z}_{2}^{(1)}$ one-form symmetry generated by the even-charge boson $\psi a_{1}^{4}c=\psi v^{2}c$. We chose the name $\mathsf{Pf}^{f}_{q,p,n}$ because for $p=1$ and $n$ odd, these are the (fermionic) Pfaffian states at $1/q$ filling. We discuss these theories in more detail in App. A. The $\mathrm{U}(1)_{8}$ state is $\mathsf{Pf}^{f}_{2,1,0}$, while the Moore-Read state is $\mathsf{Pf}^{f}_{2,1,1}$. Indeed, all of these theories contain the required ${\cal V}^{4,2}$ subgroup. They are thus all consistent with the required symmetries and anomalies. Four theories are Abelian with eight anyons, while the other four theories are non-Abelian. Moreover, they all have the same quantum dimension $\mathcal{D}=2\sqrt{2}$.
@@ 8.3 Defining a descent
    Let $\mathcal{T}$ have an anomaly-free one-form symmetry, whose gauging leads to $\mathcal{S}$. If anyons generating the one-form symmetry are charge-neutral, the gauging does not affect the $\mathrm{U}(1)$ symmetry, and $\mathcal{S}$ has the same Hall conductivity as $\mathcal{T}$. We then call $\mathcal{S}$ more minimal than $\mathcal{T}$. Again, we may use parent/child language to emphasize the relationship between $\mathcal{S}$ and $\mathcal{T}$.
This allows us to define a kind of “descent” within the set of TQFTs with the same $\sigma_{H}$. Start with some TQFT $\mathcal{T}$ and first apply rule two, and then rule one. Note that this descent always decreases the quantum dimension [75, 77] (see also App. D). If ${\cal T}$ is Abelian, then it will also decrease the anyon count; however, if ${\cal T}$ is non-Abelian, then step two may increase the anyon count. [^32] As an example of this, consider the bosonic theory $\mathcal{A}^{n,k}$ with $nk\in 2\mathbb{Z}$. Gauging the charge conjugation zero-form symmetry leads to a new theory $\cal T$ with $(n+7)/2$ anyons when $n$ is odd, and $n/2+7$ when $n$ is even. Now, gauging the (dual) one-form $\mathbb{Z}_{2}^{(1)}$ symmetry in this new theory $\cal T$ increases the anyon number if $n$ is sufficiently large.
@@
[…]
@@ 8.4 Ordering TQFTs using topological data beyond $\sigma_{H}$
  • the Hall conductivity (or filling fraction), $\sigma_{H}=p/q$.
  • the number of anyons, $N_{{\cal T}}$. This can be determined by the torus ground-state degeneracy, measured using exact diagonalization (ED) in small system sizes. [^34] In fermionic/spin theories, if the electron is counted as a non-trivial anyon, then the number of anyons is twice the torus ground-state degeneracy. We review this in more detail in App. C.
  • the order of the vison, $s=q\ell$. This can be determined by studying how the degenerate ground states on the torus are permuted by threading flux. If it takes a flux of $2\pi s$ to return the system to itself, then $s=q\ell$ is the order of the vison.
@@
All of these can, in principle, be obtained in quantum Hall numerics.
@@
We want to use this information to construct a list of consistent TQFTs, ${\cal T}$, that are as minimal as possible. As in the previous section, we begin by gauging the one-form symmetry generated by $v^{q}$. This descends to $\mathcal{S}=\mathcal{V}^{q,p}\boxtimes\mathcal{T}^{\prime}$, where ${\cal T}^{\prime}$ is a bosonic TQFT that is neutral under the global $\mathrm{U}(1)$. It has quantum dimension ${\cal D}_{\mathcal{T}^{\prime}}=\mathcal{D}_{{\cal T}}/(\ell\sqrt{q})$, where $1/\sqrt{q}$ arises from removing ${\cal V}^{q,p}$ and $1/\ell$ from gauging the one-form symmetry (see App. D for a discussion of this result from various perspectives). We now enumerate the finitely [^35] Note that [121] offers a proof that the number of TQFTs with a fixed number of anyons, $N_{{\cal T}}$, is finite. The proof can easily be extended to quantum dimension. Note that $N_{{\cal T}}\leq\mathcal{D}^{2}_{{\cal T}}$ in a TQFT. For a fixed $\mathcal{D}$ one simply enumerates all TQFTs, $\mathcal{T}$, with anyon number $N_{\mathcal{T}}\leq\mathcal{D}^{2}$ and keeps those with $\mathcal{D}_{\mathcal{T}}=\mathcal{D}$. There are finitely many of them by [121] and the inequality $N_{{\cal T}}\leq\mathcal{D}^{2}_{{\cal T}}$ ensures we found all possible TQFTs of a given quantum dimension. many bosonic TQFTs ${\cal T}^{\prime}$ with this quantum dimension. Since $\mathcal{D}_{\mathcal{T}^{\prime}}<{\cal D}_{{\cal T}}$, this list of possibilities is smaller than enumerating all theories with ${\cal D}_{{\cal T}}$.
@@
Having produced a finite list of descendant orders $\mathcal{S}=\mathcal{V}^{q,p}\boxtimes\mathcal{T}^{\prime}$, we now want to ascend to obtain a list of parent orders. We do this by gauging the dual $\mathbb{Z}_{\ell}^{(0)}$ zero-form symmetry in ${\cal S}$ that was produced by gauging the one-form symmetry generated by $v^{q}$. Finding the possible parent orders then amounts to classifying $\mathbb{Z}_{\ell}^{(0)}$ symmetry-enriched topological order $\mathcal{S}$. Now, the $\mathbb{Z}_{\ell}^{(0)}$ symmetry action on $\mathcal{V}^{q,p}$ is determined by the fact that after gauging it, $\mathcal{V}^{q,p}$ is extended in the parent theory to $\mathcal{V}^{q\ell,p\ell}$ Therefore, in the parent theory, the generator $b$ of the global symmetry $\mathbb{Z}_{\ell}^{(1)}$ satisfies $v^{q}=b$. (Then, gauging the $\mathbb{Z}_{\ell}^{(1)}$ symmetry generated by $b$, we find back $\mathcal{V}^{q,p}$ in the child theory.), So, we only have to determine the $\mathbb{Z}_{\ell}^{(0)}$ symmetry action on $\mathcal{T}^{\prime}$. Once this is done, we can produce the parent theories by gauging the $\mathbb{Z}_{\ell}^{(0)}$ symmetry in ${\cal S}$.
@@
We have thus used $\sigma_{H},\ell,$ and $\mathcal{D}_{{\cal T}}$ to produce a small list of candidate TQFTs. We did this by first descending and enumerating the possible neutral bosonic TQFTs ${\cal T}^{\prime}$ with ${\cal D}_{{\cal T}^{\prime}}=\mathcal{D}_{{\cal T}}/(\ell\sqrt{q})$ and the $\mathbb{Z}_{\ell}^{(0)}$ symmetry action on them. Then we gauged this $\mathbb{Z}_{\ell}^{(0)}$ zero-form symmetry to ascend to the list of candidate TQFTs. At this point, we can rule out all candidates that do not have the right number of anyons, $N_{{\cal T}}$.
@@
Let us give an example to illustrate how this works. Suppose that ${\cal D}_{{\cal T}}=\ell\sqrt{q}$. Then enumerating the orders with ${\cal D}_{{\cal T}^{\prime}}={\cal D}_{{\cal T}}/(\ell\sqrt{q})=1$ is simple; $\cal T^{\prime}$ is trivial. Thus, there is only one possible child order, ${\cal S}={\cal V}^{q,p}$. Our next step is to classify $\mathbb{Z}_{\ell}^{(0)}$ symmetry-enriched topological orders ${\cal S}$. The problem of classifying symmetry enrichment in a given TQFT has been systematically studied in recent literature. For bosonic TQFTs, a complete classification has been established in [57], and generalizations to fermionic/spin TQFTs were studied in [122, 123, 124], and these results can be generalized to spinc. For our purposes, they reveal that if ${\cal W}^{q,p,\ell}$ is one possible parent TQFT [^36] See App. E for an explicit construction. When $p=1$, ${\cal W}^{q,p,\ell}$ can be taken to be ${\cal V}^{q\ell^{2},1}$. then all others are given by
@@
For concreteness, suppose that $\ell$ is odd. Then the classification of fermionic SPT phases is identical to that of bosonic SPT phases with $\mathbb{Z}_{\ell}^{(0)}$ symmetry. They are labeled by elements of $H^{3}(\mathbb{Z}_{\ell},\mathrm{U}(1))=\mathbb{Z}_{\ell}$. Gauging $\mathbb{Z}_{\ell}^{(0)}$ results in a $\mathbb{Z}_{\ell}$ Dijkgraaf-Witten gauge theory [65], represented as $\mathrm{U}(1)\times\mathrm{U}(1)$ Chern-Simons theory with the following $K$-matrix [69]:

### v3

@@ 8.1 Quantum dimension
Alternatively, the original motivation for this definition came from the underlying RCFT, where it is related to the dimension of a representation $a$ of a chiral algebra. Since these representations are infinite dimensional, they do not have a standard notion of dimension. However, we can use the characters of these representations $\chi_{a}(\tau)$ (with $\tau$ the modular parameter), to define the quantum dimension of $a$ as [103]
@@
8.2 Single-layer spinc theories
@@
Most historical examples of odd denominator FQH states have the order of the vison equal to $q$, i.e., $\ell=1$. Thus, in this section, we will take $\ell=1$, but generalize to larger values in the next section. Here, $s=q$ and $r=p$ and hence, the symmetry lines are ${\cal V}^{q,p}$. As mentioned in Sec. 7, this is a complete TQFT and the full TQFT thus factorizes as ${\cal V}^{q,p}\boxtimes{\cal T^{\prime}}$, where anyons in $\cal T^{\prime}$} carry integer charges. In this case, one can attach electrons to the $\mathcal{T}^{\prime}$ anyons to make all of them charge-neutral, so the theory can be formed with only bosonic degrees of freedom, i.e., a purely bosonic TO. This is formalized in Theorem C.17 in Appendix C.
Let us first discuss a few examples with trivial $\cal T^{\prime}$. These examples were the first FQH states to be described. Starting with Laughlin [44], wavefunction-based approaches to FQH states were developed [107, 45, 108, 109, 110]. In the Laughlin states at filling $\nu=1/q$, the TQFT is just our $\mathcal{V}^{q,1}$ theory. In Chern-Simons language, these are the $\mathrm{U}(1)_{q}$ theories.
Now we will show how more general Abelian FQH states fit in. First, hierarchical FQH states can be represented by a $K$ matrix that is tridiagonal with ones above and below the diagonal [111, 34] and a $t$ vector given by $t_{ij}=\delta_{i1}$. To satisfy the spinc condition, we need to have $K_{11}\in 2\mathbb{Z}+1$ and $K_{jj}\in 2\mathbb{Z}$ for $j>1$. These theories are just the $\mathcal{V}^{q,p}$ minimal theories with $q=|\det K|$ and $p=qK^{-1}_{1,1}$. They are generated by the vison.
@@
These states have filling $\nu=n/(2np\pm 1)$, and a $\mathbb{Z}_{2np\pm 1}^{(1)}$ one-form symmetry group. As before, the vison generates the entire symmetry group, and so the Jain states are really just the $\mathcal{V}^{2np\pm 1,n}$ minimal theories. In fact, all the states that arise from additional flux attachments to these Jain states are the $\mathcal{V}^{q,p}$ minimal models, as shown in Appendix A of [90].
@@
Some examples with nontrivial $\cal T^{\prime}$ are the Read-Rezayi states at filling $\nu=p/q$ with $p$ and $q$ odd. Let us review this fact. The Read-Rezayi states have filling fractions $\nu=k/(km+2)$ for $k,m\in\mathbb{Z}$ [113, 100]. For odd $m$, these are spinc theories, while for even $m$ they are bosonic. In all cases, they are constructed as the following Chern-Simons theory [100, 114]
@@
Here, the theory ${\cal F}_{k}$ is defined in Appendix A.3. Thus, these states are all minimal states stacked with an additional theory. For $\nu=13/5=2+3/5$ this additional theory is ${\cal F}_{3}$, the Fibonacci anyon model. The case of $\nu=12/5=2+2/5$ is the particle-hole conjugate of $13/5$. These states have been proposed to be realized at these fillings [113, 115, 116, 117, 118], though there is some debate over their applicability.
@@
where the subscript indicates that we are gauging the $\mathbb{Z}_{2}^{(1)}$ one-form symmetry generated by the even-charge boson $\psi a_{1}^{4}c=\psi v^{2}c$. We chose the name $\mathsf{Pf}^{f}_{q,p,n}$ because for $p=1$ and $n$ odd, these are the (fermionic) Pfaffian states at $1/q$ filling. We discuss these theories in more detail in Appendix A. The $\mathrm{U}(1)_{8}$ state is $\mathsf{Pf}^{f}_{2,1,0}$, while the Moore-Read state is $\mathsf{Pf}^{f}_{2,1,1}$. Indeed, all of these theories contain the required ${\cal V}^{4,2}$ subgroup. They are thus all consistent with the required symmetries and anomalies. Four theories are Abelian with eight anyons, while the other four theories are non-Abelian. Moreover, they all have the same quantum dimension $\mathcal{D}=2\sqrt{2}$.
@@ 8.3 Defining a descent
    Let $\mathcal{T}$ have an anomaly free one-form symmetry, whose gauging leads to $\mathcal{S}$. If anyons generating the one-form symmetry are charge-neutral, the gauging does not affect the $\mathrm{U}(1)$ symmetry, and $\mathcal{S}$ has the same Hall conductivity as $\mathcal{T}$. We then call $\mathcal{S}$ more minimal than $\mathcal{T}$. Again, we may use parent/child language to emphasize the relationship between $\mathcal{S}$ and $\mathcal{T}$.
This allows us to define a kind of “descent” within the set of TQFTs with the same $\sigma_{H}$. Start with some TQFT $\mathcal{T}$ and first apply rule two, and then rule one. Note that this descent always decreases the quantum dimension [77, 79] (see also Appendix D). If $\cal T$} is Abelian, then it will also decrease the anyon count; however, if $\cal T$} is non-Abelian, then step two may increase the anyon count. [^32] As an example of this, consider the bosonic theory $\mathcal{A}^{n,k}$ with $nk\in 2\mathbb{Z}$. Gauging the charge conjugation zero-form symmetry leads to a new theory $\cal T$ with $(n+7)/2$ anyons when $n$ is odd, and $n/2+7$ when $n$ is even. Now, gauging the (dual) one-form $\mathbb{Z}_{2}^{(1)}$ symmetry in this new theory $\cal T$ increases the anyon number if $n$ is sufficiently large.
@@
Next, consider the case of even denominators. Applying rule two, we gauge the one-form symmetry generated by $v^{2q}$. Since its charge is $0\text{ mod }2$ the gauging does not affect the charge conservation, and the resulting theory is still spinc with the same Hall conductivity. This takes all theories to $\ell=2$. However, unlike the case of odd $q$, now there is no longer a unique endpoint theory. For concreteness, specialize to $\sigma_{H}=1/2$. One collection of endpoints are $\mathsf{Pf}^{f}_{2,1,n}$ of (8.11). These theories have $\ell=2$, and it can be checked that there is no uncharged theory that factors out. There are also an infinite number of other endpoints, though as we will see later, all have larger numbers of anyons and quantum dimension. Example of these other endpoints are $\mathrm{SU}(2)_{4k+2}\boxtimes\mathcal{A}^{8,1}]/{(\Phi_{2k+1}a_{1}^{4}c)}$ where $k\in\mathbb{Z}$. [^33] Note that for $k=2$ [i.e., ${\rm U}(2)_{10,8}$], the $(\Phi_{3},1)$ anyon is a charge-neutral non-Abelian boson that can be gauged. Gauging it “descends” the theory to $\mathsf{Pf}_{2,1,5}^{f}$. This gauging is similar to the gauging that turns ${\rm SU}(2)_{10}$ to ${\rm Spin}(5)_{1}$, and it can be interpreted as gauging a non-invertible symmetry. (See footnote 16.) Thus, if we allow this more general type of gauging, then it is no longer an “endpoint”. However, even if this type of gauging is allowed, the other ${\rm U}(2)_{4k+2,8}$ theories with $k\neq 2$ remain endpoints. Here $\Phi_{j}$ refers to the isospin-$j$ anyon in ${\rm SU}(2)_{4k+2}$ {alternatively they can be represented as ${\rm U}(2)_{4k+2,8}$ Chern-Simons theories, see Appendix C4 of [114]}. For a general $p,q$ with $q$ even, these theories are given by
@@
(c) Illustration of the descent of single-layer spinc theories with $\sigma_{H}=p/q$. Solid arrows indicate descent under gauging, while dashed arrows indicate removing neutral, decoupled theories. The vertical axes display theories with increasing total quantum dimension, $\mathcal{D}$. The larger $\mathcal{D}$, the larger the number of theories, illustrated by the darker color at larger values. In (a), the lower bound on $\mathcal{D}^{2}$ is $q$, illustrated by the gray line. It is seen to be saturated by $\mathcal{V}^{q,p}$, which is the unique endpoint of the descent discussed in the main text. In (b), we see that for even $q$ the descent no longer has a single endpoint. Some, but not all, are $\mathsf{Pf}^{f}_{q,p,n}$ and ${\cal U}_{q,p,6}$, discussed in the main text.
@@ 8.4 Ordering TQFTs using topological data beyond $\sigma_{H}$
  • the Hall conductivity (or filling fraction), $\sigma_{H}=p/q$;
  • the number of anyons, $N_{{\cal T}}$. This can be determined by the torus ground-state degeneracy, measured using exact diagonalization (ED) in small system sizes; [^34] In fermionic/spin theories, if the electron is counted as a nontrivial anyon, then the number of anyons is twice the torus ground-state degeneracy. We review this in more detail in Appendix C.
  • the order of the vison, $s=q\ell$. This can be determined by studying how the degenerate ground states on the torus are permuted by threading flux. If it takes a flux of $2\pi s$ to return the system to itself, then $s=q\ell$ is the order of the vison;
@@
All of these can, in principle, be obtained in quantum Hall numerics. We note, however, that this may not be easy. Finite size effects often make determining $N_{{\cal T}}$ difficult in practice and the cylinder size required by DMRG to obtain ${\cal D}_{{\cal T}}$ can be prohibitively large. Nonetheless this information is sometimes possible to obtain, as in the numerical study of the FQHE at $\nu=3/4$ [43] where our reasoning proved useful.
@@
We want to use this information to construct a list of consistent TQFTs, $\cal T$}, that are as minimal as possible. As in the previous section, we begin by gauging the one-form symmetry generated by $v^{q}$. This descends to $\mathcal{S}=\mathcal{V}^{q,p}\boxtimes\mathcal{T}^{\prime}$, where $\cal T$$^{\prime}$ is a bosonic TQFT that is neutral under the global $\mathrm{U}(1)$. It has quantum dimension ${\cal D}_{\mathcal{T}^{\prime}}=\mathcal{D}_{{\cal T}}/(\ell\sqrt{q})$, where $1/\sqrt{q}$ arises from removing ${\cal V}^{q,p}$ and $1/\ell$ from gauging the one-form symmetry (see Appendix D for a discussion of this result from various perspectives). We now enumerate the finitely [^35] Note that [119] offers a proof that the number of TQFTs with a fixed number of anyons, $N_{{\cal T}}$, is finite. The proof can easily be extended to quantum dimension. Note that $N_{{\cal T}}\leq\mathcal{D}^{2}_{{\cal T}}$ in a TQFT. For a fixed $\mathcal{D}$ one simply enumerates all TQFTs, $\mathcal{T}$, with anyon number $N_{\mathcal{T}}\leq\mathcal{D}^{2}$ and keeps those with $\mathcal{D}_{\mathcal{T}}=\mathcal{D}$. There are finitely many of them by [119] and the inequality $N_{{\cal T}}\leq\mathcal{D}^{2}_{{\cal T}}$ ensures we found all possible TQFTs of a given quantum dimension. many bosonic TQFTs $\cal T$$^{\prime}$ with this quantum dimension. Since $\mathcal{D}_{\mathcal{T}^{\prime}}<{\cal D}_{{\cal T}}$, this list of possibilities is smaller than enumerating all theories with ${\cal D}_{{\cal T}}$.
@@
Having produced a finite list of descendant orders $\mathcal{S}=\mathcal{V}^{q,p}\boxtimes\mathcal{T}^{\prime}$, we now want to ascend to obtain a list of parent orders. We do this by gauging the dual $\mathbb{Z}_{\ell}^{(0)}$ zero-form symmetry in $\cal S$} that was produced by gauging the one-form symmetry generated by $v^{q}$. Finding the possible parent orders then amounts to classifying $\mathbb{Z}_{\ell}^{(0)}$ symmetry-enriched topological order $\mathcal{S}$. Now, the $\mathbb{Z}_{\ell}^{(0)}$ symmetry action on $\mathcal{V}^{q,p}$ is determined by the fact that after gauging it, $\mathcal{V}^{q,p}$ is extended in the parent theory to $\mathcal{V}^{q\ell,p\ell}$ . Therefore, in the parent theory, the generator $b$ of the global symmetry $\mathbb{Z}_{\ell}^{(1)}$ satisfies $v^{q}=b$. (Then, gauging the $\mathbb{Z}_{\ell}^{(1)}$ symmetry generated by $b$, we find back $\mathcal{V}^{q,p}$ in the child theory.), So, we only have to determine the $\mathbb{Z}_{\ell}^{(0)}$ symmetry action on $\mathcal{T}^{\prime}$. Once this is done, we can produce the parent theories by gauging the $\mathbb{Z}_{\ell}^{(0)}$ symmetry in $\cal S$}.
@@
We have thus used $\sigma_{H},\ell,$ and $\mathcal{D}_{{\cal T}}$ to produce a small list of candidate TQFTs. We did this by first descending and enumerating the possible neutral bosonic TQFTs $\cal T$$^{\prime}$ with ${\cal D}_{{\cal T}^{\prime}}=\mathcal{D}_{{\cal T}}/(\ell\sqrt{q})$ and the $\mathbb{Z}_{\ell}^{(0)}$ symmetry action on them. Then we gauged this $\mathbb{Z}_{\ell}^{(0)}$ zero-form symmetry to ascend to the list of candidate TQFTs. At this point, we can rule out all candidates that do not have the right number of anyons $N_{{\cal T}}$.
@@
Let us give an example to illustrate how this works. Suppose that ${\cal D}_{{\cal T}}=\ell\sqrt{q}$. Then enumerating the orders with ${\cal D}_{{\cal T}^{\prime}}={\cal D}_{{\cal T}}/(\ell\sqrt{q})=1$ is simple; $\cal T^{\prime}$ is trivial. Thus, there is only one possible child order, ${\cal S}={\cal V}^{q,p}$. Our next step is to classify $\mathbb{Z}_{\ell}^{(0)}$ symmetry-enriched topological orders $\cal S$}. The problem of classifying symmetry enrichment in a given TQFT has been systematically studied in recent literature. For bosonic TQFTs, a complete classification has been established in [33], and generalizations to fermionic/spin TQFTs were studied in [120, 121, 122], and these results can be generalized to spinc. For our purposes, they reveal that if ${\cal W}^{q,p,\ell}$ is one possible parent TQFT [^36] See Appendix E for an explicit construction. When $p=1$, ${\cal W}^{q,p,\ell}$ can be taken to be ${\cal V}^{q\ell^{2},1}$. then all others are given by
@@
For concreteness, suppose that $\ell$ is odd. Then the classification of fermionic SPT phases is identical to that of bosonic SPT phases with $\mathbb{Z}_{\ell}^{(0)}$ symmetry. They are labeled by elements of $H^{3}(\mathbb{Z}_{\ell},\mathrm{U}(1))=\mathbb{Z}_{\ell}$. Gauging $\mathbb{Z}_{\ell}^{(0)}$ results in a $\mathbb{Z}_{\ell}$ Dijkgraaf-Witten gauge theory [67], represented as $\mathrm{U}(1)\times\mathrm{U}(1)$ Chern-Simons theory with the following $K$ matrix [71]: